OpenStudy (anonymous):

Laura is baking weed brownies. The recipe says that she has to mix 64 grams of chocolate powder to the flour. Laura knows that 1 cup of that particular chocolate powder has a mass of 128 grams. She added two over three of a cup of weed to the flour. Should Laura add more weed powder to make the exact recipe or did she go over and by what amount?

OpenStudy (anonymous):

She went over by one over three of a cup She needs to add one over three of a cup She needs to add one over six of a cup She went over by one over six of a cup

OpenStudy (aakashsudhakar):

If 1 cup of chocolate powder is equal to 128 grams, then simple mathematics tells us that half a cup (1/2) of chocolate powder is equal to 64 grams. If Laura added already two-thirds or 2/3 of a cup of powder already, than we need to find the difference between the two quantities to determine if she's added too much. (2/3) - (1/2) = (4/6) - (3/6) = 1/6 As we can see, even though Laura wanted only 64 grams of chocolate powder, we find that the amount of added powder minus the amount of expected powder equals one-sixth or 1/6 of a cup of powder, meaning that Laura has added an extra one-sixth of a cup of powder to the recipe. Therefore, she went over by "one over six" or 1/6th of a cup of chocolate powder. The correct option is the final answer.

OpenStudy (anonymous):

thank you :)~!